# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix F Using the Performance Difference Lemma
where ${\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]})=\widehat{\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{u},F_{s_{\Delta\setminus u}})$ is the global agent’s action and ${\pi}_{k,m}^{i,\mathrm{est}}(s_{g},s_{[n]}):=\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Here, $\Delta_{i}$ is a random variable supported on (${[n]\setminus i\choose k-1}$, $u$ is a random variable uniformly distributed on $[n]$, and $\Delta$ is a random variable uniformly distributed on $[n]\setminus u$. Then, denote the optimal policy $\pi^{*}$ given by
@@
where $\pi^{*}_{g}(s_{g},s_{[n]})$ is the global agent’s action, and $\pi_{i}^{*}(s_{g},s_{[n]}):=\pi^{*}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Next, to compare the difference in the performance of $\pi^{*}(s)$ and ${\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{[n]})$, we define the value function of a policy $\pi$ to be the infinite-horizon $\gamma$-discounted rewards, (denoted $V^{\pi}$) as follows:
Definition F.2.
@@
$V^{\pi^{*}}(s_{0})-V^{{\pi}^{\mathrm{est}}_{k,m}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}^{\mathrm{est}}_{k,m}}(s_{0})$ $=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}[V^{\pi^{*}}(s)-Q^{\pi^{*}}(s,a)]$
@@
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}}\bigg{[}Q^{\pi^{*}}(s,\pi^{*}(\cdot|s))-\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}Q^{\pi^{*}}(s,a)\bigg{]}$
@@
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\sum_{\Delta^{2}\in{[n]\setminus 2\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
@@
$-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))\bigg{)}$
@@
$V^{\pi^{*}}(s_{0})$ $-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$
$\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim\widetilde{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{i}\in{[n]\setminus i\choose k-1},\\ \forall i\in[n]\end{subarray}}\frac{1}{n{n\choose k}{n-1\choose k-1}^{n}}\bigg{(}\sum_{i\in[n]}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$+\frac{\frac{1}{n}}{{n\choose k}{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}\bigg{)}$
Lemma F.4 shows a uniform bound on
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
(independent of $\Delta^{i},\forall i\in[n]$), allowing the sums and the counts in the denominator will cancel out.
Observe that $\widehat{\pi}_{k,m}^{*}:\mathcal{S}\to\mathcal{A}$ and $\pi^{*}:\mathcal{S}\to\mathcal{A}$ are deterministic functions. Therefore, denote $a=\pi^{*}(s)$. Then, from Lemma F.6, we have that
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},\widehat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{g},\widehat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{1:n})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Similarly, from Corollary F.6.1, we have that
$\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\bigg{|}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{2\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
which yields the claim. We defer parameter optimization to Lemma F.7.∎
Lemma F.4 (Uniform Bound on $Q^{*}$ with different actions).
For all $s\in S,\Delta\in{[n]\choose k}$ and $\Delta^{i}\in{[n]\setminus i\choose k-1}$,
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$+\left|\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\right|$

### v3

@@ Appendix G Using the Performance Difference Lemma
where ${\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]})=\hat{\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{u},F_{s_{\Delta\setminus u}})$ is the global agent’s action and ${\pi}_{k,m}^{i,\mathrm{est}}(s_{g},s_{[n]}):=\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Here, $\Delta_{i}$ is a random variable supported on (${[n]\setminus i\choose k-1}$, $u$ is a random variable uniformly distributed on $[n]$, and $\Delta$ is a random variable uniformly distributed on $[n]\setminus u$. Then, denote the optimal policy $\pi^{*}$ given by
@@
where $\pi^{*}_{g}(s_{g},s_{[n]})$ is the global agent’s action, and $\pi_{i}^{*}(s_{g},s_{[n]}):=\pi^{*}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Next, in order to compare the difference in the performance of $\pi^{*}(s)$ and ${\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{[n]})$, we define the value function of a policy $\pi$ to be the infinite-horizon $\gamma$-discounted rewards, (denoted by $V^{\pi}$) as follows:
Definition G.2.
@@
$V^{\pi^{*}}(s_{0})-V^{{\pi}^{\mathrm{est}}_{k,m}}(s_{0})\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\tilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}^{\mathrm{est}}_{k,m}}(s_{0})$ $=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}[V^{\pi^{*}}(s)-Q^{\pi^{*}}(s,a)]$
@@
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\tilde{\pi}_{k,m}^{\mathrm{est}}}}\bigg{[}Q^{\pi^{*}}(s,\pi^{*}(\cdot|s))-\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}Q^{\pi^{*}}(s,a)\bigg{]}$
@@
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{1},\dots,\Delta^{n}:\\ \Delta^{i}\in{[n]\setminus i\choose k-1}\end{subarray}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
@@
$-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))\bigg{)}$
@@
$V^{\pi^{*}}(s_{0})$ $-V^{\tilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$
$\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\tilde{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim\tilde{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{i}\in{[n]\setminus i\choose k-1},\\ \forall i\in[n]\end{subarray}}\frac{1}{n{n\choose k}{n-1\choose k-1}^{n}}\bigg{(}\sum_{i\in[n]}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$+\frac{\frac{1}{n}}{{n\choose k}{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}\bigg{)}$
Lemma G.4 shows a uniform bound on
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
(independent of $\Delta^{i},\forall i\in[n]$), allowing the sums and the counts in the denominator will cancel out. Observe that $\hat{\pi}_{k,m}^{*}:\mathcal{S}\to\mathcal{A}$ and $\pi^{*}:\mathcal{S}\to\mathcal{A}$ are deterministic functions. Therefore, denote $a=\pi^{*}(s)$. Then, from Lemma G.6,
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{1},\dots,\Delta^{n}:\\ \Delta^{i}\in{[n]\setminus i\choose k-1}\end{subarray}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\hat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},\hat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{g},\hat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{1:n})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Similarly, from Corollary G.7, we have that
$\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\bigg{|}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},$
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})-Q^{*}(s,\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{2\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\tilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
which yields the claim. We defer parameter optimization to Lemma G.8.∎
Lemma G.4 (Uniform Bound on $Q^{*}$ with different actions).
For all $s\in\mathcal{S},\Delta\in{[n]\choose k}$ and $\Delta^{i}\in{[n]\setminus i\choose k-1}$ for $i\in[n]$, we have:
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\quad\quad\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad\quad\quad\quad+\bigg{|}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$